# Tap into Reality: Understanding the Impact of Interactions on Presence and Reaction Time in Mixed Reality

### v1

@@ 3.6 Procedure
In within-subjects designs, managing order effects is essential to measure the experimental conditions’ impact accurately. To handle this, we used a balanced Latin square design [47] to randomize the order of the four conditions across the participants. Each condition was placed in each position an equal number of times and directly followed others equally. The 4x4 Latin square pattern, repeatedly multiple times in a round-robin manner, enabled proper distribution: {DM, SM, DS, SS }, { SM, DS, SS, DM }, {DS, SS, DM, SM }, {SS, DM, SM, DS } This approach made sure that each condition occupied every position once per set of four participants. In all conditions, participants wore the headset and performed two successful interactions (bubble popping and box moving) to familiarize themselves before each condition. These practice interactions were recorded but excluded from the analysis. As shown in Figure 6, periodic cues are generated (3s in Select and 5s in Manipulate), and participants answered a prompt question halfway through each condition (1.5 minutes in). Participants removed the headset after three minutes, completing all trials (35 in Manipulate conditions and 20 in Select conditions) and answering the prompt question. After each condition, they filled out presence questionnaires and responded to an open-ended question. After completing all four conditions, we conducted a small debriefing session where we answered their questions. The study session, including the initial briefing, four conditions, questionnaires, and debrief, took about 45 minutes.
@@ 4.1 Statistical Analysis
We used Levene’s homogeneity test, which held true for all dependent variables. The Shapiro-Wilk test found that the aggregate questionnaire, individual questionnaires, prompt score, and reaction time were normally distributed. Among the different subscales, self-evaluation of performance (EVAL), possibility to examine (EXAM), and general presence (GP) failed the normality test, while other subscales were found to be normally distributed. We do not evaluate the impact of our dependent variables on the three aforementioned subscales, in addition to spatial presence (SP), as summary statistics show that interaction type and task type did not significantly impact these subscales.
@@
Table 3 presents the results of the two-way repeated measures ANOVA, using combined questionnaire scores, individual subscores, prompt scores, and reaction time as the dependent variables. The analysis reveals that the main effect of interaction type is statistically significant across several subscales and dependent variables. For the ACT subscale, the interaction type shows the strongest effect, with $F=78.79,p<0.05,\eta^{2}=0.276$, indicating that interaction type explains 27.6% of the variance in this subscale. This finding highlights that interaction (D vs. S) substantially influences participants’ perception of their ability to act within the environment. The IFQUAL subscale shows a significant main effect of interaction type, with $F=36.72,p<0.05,\eta^{2}=0.152$, indicating 15.2% of the variance. This suggests that the quality of the interaction interface is influenced by whether the interaction is D and S. For PQ-REAL, the interaction type effect is significant, with $F=53.85,p<0.05,\eta^{2}=0.203$, indicating that interaction type accounts for 20.3% of the variance in the perceived realism of the VE. In INV, the interaction type effect is significant, with $F=31.93,p<0.05,\eta^{2}=0.128$, showing that interaction type explains 12.8% of the variance in involvement scores. The IPQ-REAL subscale also exhibits a significant main effect of interaction type, with $F=22.00,p<0.05,\eta^{2}=0.098$, meaning that interaction type indicates 9.8% of the variance in this measure of realism. For the prompt score, which captures participants’ real-time sense of presence, the effect of interaction type is marginally significant, with $F=3.899,p=0.0497,\eta^{2}=0.019$, explaining 1.9% of the variance. Regarding reaction time, the interaction type exerts a significant effect, with $F=54.43,p<0.05,\eta^{2}=0.163$. This shows that the type of interaction significantly influences how quickly participants respond during the tasks.
@@ 4.1.1 Task Type Effects
The main effect of task type is statistically significant for several subscales. For the Possibility to Act (ACT) subscale, task type shows an effect with $F=9.927,p<0.05,\eta^{2}=0.035$. This indicates that the type of task (Select vs. Manipulate) influences participants’ perception of their ability to act in the virtual environment. For IFQUAL, task type shows a significant effect, with $F=8.794,p<0.05,\eta^{2}=0.036$. This suggests that task type impacts how participants evaluate the quality of the interface. For PQ-REAL, task type shows a significant effect with $F=15.98,p<0.05,\eta^{2}=0.060$, 6% of the variance in realism scores. The INV exhibits a significant task type effect, with $F=21.37,p<0.05,\eta^{2}=0.085$, indicating that task type influences participants’ level of involvement in the environment. For IPQ-REAL, task type shows a significant effect, with $F=7.307,p<0.05,\eta^{2}=0.032$. In terms of the prompt score, task type exerts a small but statistically significant effect, with $F=4.591,p<0.05,\eta^{2}=0.022$, indicating 2.2% of the variance. For reaction time, task type has the most substantial effect, with $F=74.62,p<0.05,\eta^{2}=0.223$, showing that the type of task significantly influences how quickly participants respond.
@@ 4.1.3 Correlation
In Figure 8, the relationship between presence scores and reaction time is visualized, with each plot representing one of the four experimental conditions. Pearson’s correlation coefficients across conditions showed a modest negative correlation between presence and reaction time, with values ranging from -0.42 in DM to -0.63 in SM with linear regression models fit, illustrating the inverse relationship between presence and reaction time across conditions.
@@ 4.2 Mediation Analysis
We conducted the mediation analysis ( Figure 7) to understand the relationship between interaction, task, and reaction time, with presence as a mediator, by using a bootstrapping approach with 1000 resamples with 95% confidence intervals (CI). The analysis revealed a positive effect of interaction type on presence, with a coefficient of $0.475$ ($p<0.05$, $CI[0.380,0.573]$). This suggests that D resulted in higher levels of presence than S interactions, indicating a strong relationship between interaction type and presence. For the task type, we also examined its effect on presence. The results showed that Select generally resulted in higher levels of presence than Manipulate tasks. For Task Type, the effect on presence was slightly smaller, with a coefficient of $0.362$ ($p<0.05$, $CI[0.245,0.485]$), suggesting that Select tasks led to higher levels of presence than Manipulate tasks. The next phase of our analysis focused on the relationship between presence and reaction time. We found that Presence significantly negatively affected reaction time with a coefficient of $-0.610(p<0.05$, $CI[0.245,0.485]$), suggesting that as presence increased, reaction time became faster (also in  Figure 8).
@@
The indirect effect of interaction type on reaction time, mediated by Presence, was calculated to be $-0.290(p<0.05$, $CI[-0.398,-0.190]$). This suggests that the improvement in reaction time associated with interaction type was largely mediated by the sense of presence. We also calculated the indirect effect of task type on reaction time, again mediated by Presence. This indirect effect was $-0.221(p<0.05$, $CI[-0.310,-0.130]$), showing that Select improved presence, leading to faster reaction times. The direct effect of interaction type on reaction time, controlling for presence, is found to be $0.017(p>0.05$, $CI[-0.087,0.122]$) 0.017, which is small and not statistically significant. This indicates that once presence is accounted for, interaction type had little direct influence on reaction time. This suggests that the effect of interaction type on reaction time is primarily mediated by presence. The direct effect of task type on reaction time, controlling for presence, is $0.092(p>0.05$, $CI[-0.015,0.199]$), indicating a small positive effect of Select on reaction time, meaning that this direct effect is not statistically significant. Thus, the relationship between task type and reaction time is explained through the presence of the mediator.
@@ 4.4 Interpretation of Results
In this section, we interpret the effects identified in our statistical analysis and explore the reasons for their manifestation. Our study used the overall change in presence as the interaction mechanism changed as the first check; the relationship between presence and reaction time would not be observed if the presence did not change. We observed a significant difference in overall presence, which is found to be statistically significant (Table 3). Therefore, we accept H1. We also found that the change in the interaction mechanism changed the overall presence, and the individual subscales showed consistent statistical significance. The interaction mechanisms differed in enabling users to act, the interface quality, how realistic they felt, and the level of involvement they needed to work. As a result, the interaction type is the dominant independent factor that significantly impacted the presence score across subscales. Furthermore, as the direct interactions were more involved than the S, the presence score for DS and DM is higher (Table 2) and statistically significant, and mediation analysis also supports H1.1 as the indirect effect of interaction type on reaction time, mediated by presence, shows that D result in higher presence. Thus, H1.1 can be accepted.
The different tasks differed in the level of involvement they expected from the participant; it is expected that changing a task would impact presence. However, we also expected that the task’s nature might impact the interaction type’s effect on presence; for example, the participant may find a given interaction mechanism better suited to one task than the other. However, our results suggested that while task impacts presence, the effect of the two interaction types on presence does not change. We did not observe an interaction effect at the overall questionnaire and subscales levels (omitted for brevity). Therefore, H2 is rejected. Our next hypothesis took a more specific approach to the cross-over between task and interaction, stating that there will be a change in presence if the interaction changes. This is observed, and H2.1 can be accepted.
@@ 4.5 Thematic Analysis
Involvement: As both quantitative and qualitative results suggested, transitioning between different experimental conditions often caused participants to feel a heightened sense of involvement. Comments such as “Being able to control the VE directly was engrossing” and “I felt really part of the virtual world when the interactions were responsive” were common in DM and a few in DS.
Realism: Similarly, the shift to conditions DM and DS was noted by the participants who commented, “..objects looked almost tangible,” and ”The realistic response of the environment to my actions added to the fun”. This suggests that enhancements in visual and interaction realism contribute to the sense of presence.
Interface Quality: Feedback on interface quality often highlighted the limitations of the field of view and control responsiveness. For example, one participant noted, “..small field of view limited my interaction,” while another said, “Sometimes the gestures did not register, which pulled me out of the experience” (reported in Select). These comments align with observations where interface limitations negatively impact user experience.
Comfort: The physical comfort of using the headset also played a crucial role in the presence experience. Participants frequently mentioned comfort issues, stating, “The headset felt heavy after a while,” or ”I experienced some eye strain during the last few trials.”
@@ 5 Discussion
Our findings revealed significant effects of interaction and task types on presence scores, as indicated by our two-way repeated measures ANOVA. The interaction type showed a higher influence on the presence scores across all subscales, while task type had a more pronounced effect on reaction time. Notably, the interaction between task and interaction type had a negligible impact, suggesting that the influence of interaction type on presence and reaction time is robust across different tasks.
@@
Correlation between Presence and Reaction Time. The correlation between presence and reaction time was modestly negative, ranging from -0.42 to -0.63 across different scenarios. This correlation is visualized in Figure 8, with linear regression lines illustrating the relationship between these measures. The trend indicates that as presence increases, reaction time decreases, suggesting that higher levels of presence in an MR can lead to faster responses.
@@
Interpretation of Hypotheses. H1 and H2 were supported, indicating that direct interactions, which are more immersive, consistently enhance the sense of presence compared to symbolic interactions. This is reflective of the intuitive nature of direct interactions that closely mimic real-world actions, facilitating a stronger engagement with the VE. H3 was not supported, which highlights that the superior efficacy of direct interaction in enhancing presence is not contingent on the task type. This suggests a general preference or suitability of direct interactions over symbolic ones in MR, regardless of the specific task being performed. H2.1 was confirmed, underscoring that changes in interaction type are sufficient to alter the sense of presence significantly. This finding is crucial for designers of MR interfaces, as it emphasizes the importance of choosing appropriate interactions to enhance the overall experience. H3, H3.1, and H3.2 highlight the relationships between interaction types, reaction times, and presence. These results confirm that interaction types not only affect presence but also significantly impact task performance efficiency, as reflected in reaction times.
@@
Theoretical and Practical Implications. The implications of these results are multifaceted. Theoretically, they contribute to the understanding of how different interactions influence user experience in MR, offering insights into the psychological and physiological aspects of presence. Practically, these findings can guide the development of more effective MR applications, where the choice of interaction type can be tailored to improve both the sense of presence and the efficiency of performance (reaction time).
@@ 6 Limitations and Future Work
In this section, we briefly discuss the limitations of using reaction time as a presence measure as well as our study. We also outline the potential future directions for this work.
One challenge we faced was the use of presence questionnaires originally designed for VR environments in an MR context. Due to the lack of validated MR-specific presence measures, we opted to use VR questionnaires, as they offer the most established and reliable way to assess presence. Additionally, the questionnaire score can be used to indicate the level of presence a user is experiencing. However, since reaction time also depends on the task at hand, just looking at it does not provide enough information about the level of presence. At present, reaction time is best suited for measuring the change in presence rather than the absolute level of presence. Further studies should explore whether the effect of task type on reaction time can be modeled, as this may enable the use of reaction time as a standalone absolute measure of presence. We focused on two interaction types (D and S) and two tasks (Select and Manipulate) for clarity and control in our study, allowing for a focused examination of foundational MR interactions before exploring more complex variations in future research. Lastly, further research could explore the long-term effects of these interactions on user adaptation and learning. Analyzing prolonged exposure to varied interactions will reveal how users adapt over time, potentially developing new strategies or behaviors. This research could uncover patterns in the correlation between sustained exposure and presence scores and how extended exposure influences reaction times. Exploring multisensory feedback integration can significantly enhance immersion and presence. Researchers could examine how these additional sensory channels influence the relationship between presence and reaction time by incorporating auditory, haptic, and olfactory cues. VE that allow multi-user interactions provide a unique opportunity to study social presence. Researching how users perceive others in shared virtual spaces and how collaboration impacts interactions, overall presence, and reaction time could be an exciting avenue.

### v2

@@ 3.6 Procedure
In within-subjects designs, managing order effects is essential to measure the experimental conditions’ impact accurately. To handle this, we used a balanced Latin square design [49] to randomize the order of the four conditions across the participants. Each condition was placed in each position an equal number of times and directly followed others equally. The 4x4 Latin square pattern, repeatedly multiple times in a round-robin manner, enabled proper distribution: {DM, SM, DS, SS }, { SM, DS, SS, DM }, {DS, SS, DM, SM }, {SS, DM, SM, DS } This approach made sure that each condition occupied every position once per set of four participants. In all conditions, participants wore the headset and performed two successful interactions (bubble popping and box moving) to familiarize themselves before each condition. These practice interactions were recorded but excluded from the analysis. As shown in Figure 6, periodic cues are generated (3s in Select and 5s in Manipulate), and participants answered a prompt question halfway through each condition (1.5 minutes in). Participants removed the headset after three minutes, completing all trials (35 in Manipulate conditions and 20 in Select conditions) and answering the prompt question. After each condition, they filled out presence questionnaires and responded to an open-ended question (general comments used for qualitative analysis in §4.5). After completing all four conditions, we conducted a small debriefing session where we answered their questions. The study session, including the initial briefing, four conditions, questionnaires, and debrief, took about 45 minutes.
@@ 4.1 Statistical Analysis
We used Levene’s homogeneity test, which held true for all dependent variables. The Shapiro-Wilk test found that the aggregate questionnaire, individual questionnaires, prompt score, and reaction time were normally distributed. Among the different subscales, self-evaluation of performance (EVAL), possibility to examine (EXAM), and general presence (GP) failed the normality test, while other subscales were found to be normally distributed. We do not evaluate the impact of our dependent variables on the three aforementioned subscales, in addition to spatial presence (SP), as summary statistics show that interactions and tasks did not significantly impact these subscales.
@@
Table 3 presents the results of the two-way repeated measures ANOVA, using combined questionnaire scores, individual subscores, prompt scores, and reaction time as the dependent variables. The analysis reveals that the main effect of interaction type is statistically significant across several subscales and dependent variables. The interaction type shows a significant effect within the study’s constraints ($F=78.79,p<0.05,\eta^{2}=0.276$), indicating that interaction type explains 27.6% variance in ACT subscale for the tested conditions. This finding highlights that interaction (D vs. S) substantially influences participants’ perception of their ability to act within the environment. The IFQUAL subscale shows a significant main effect of interaction type ($F=36.72,p<0.05,\eta^{2}=0.152$), indicating 15.2% variance. This suggests that the quality of the interaction interface is influenced by whether the interaction is D and S. For PQ-REAL, the interaction type effect is significant ($F=53.85,p<0.05,\eta^{2}=0.203$), indicating that interaction type accounts for 20.3% of the variance in the perceived realism of the VE. In INV, the interaction type effect is significant, with $F=31.93,p<0.05,\eta^{2}=0.128$, showing that interaction type explains 12.8% of the variance in involvement scores. The IPQ-REAL subscale also exhibits a significant main effect of interaction type ($F=22.00,p<0.05,\eta^{2}=0.098$), meaning that interaction type indicates 9.8% of the variance in this measure of realism. For the prompt score, which captures participants’ real-time sense of presence, the effect of interaction type is marginally significant, with $F=3.899,p=0.0497,\eta^{2}=0.019$, explaining 1.9% of the variance. The interaction type significantly affects reaction time ($F=54.43,p<0.05,\eta^{2}=0.163$), influencing participant reaction time in tested conditions.
@@ 4.1.1 Task Type Effects
The main effect of task type is statistically significant for several subscales. For the Possibility to Act (ACT) subscale, task type shows an effect with $F=9.927,p<0.05,\eta^{2}=0.035$. This indicates that the type of task (Select vs. Manipulate) influences participants’ perception of their ability to act in the virtual environment. For IFQUAL, task type shows a significant effect, with $F=8.794,p<0.05,\eta^{2}=0.036$. This suggests that task type impacts how participants evaluate the quality of the interface. For PQ-REAL, task type shows a significant effect with $F=15.98,p<0.05,\eta^{2}=0.060$, 6% of the realism scores variance. The INV exhibits a significant task type effect, with $F=21.37,p<0.05,\eta^{2}=0.085$, indicating that task type influences participants’ level of involvement in the environment. For IPQ-REAL, task type shows a significant effect, with $F=7.307,p<0.05,\eta^{2}=0.032$. Regarding the prompt score, task type exerts a small but statistically significant effect, with $F=4.591,p<0.05,\eta^{2}=0.022$, indicating 2.2% variance. For reaction time, task type has the most substantial effect, with $F=74.62,p<0.05,\eta^{2}=0.223$, showing that the type of task significantly influences how quickly participants respond.
@@ 4.1.3 Correlation
In Figure 8, the relationship between presence scores and reaction time is visualized, with each plot representing one of the four experimental conditions. Pearson’s correlation coefficients across tested conditions showed a modest negative correlation between presence and reaction time, supporting the relationship between interaction type, task type, and user performance in the context of M and S tasks, with values ranging from -0.42 in DM to -0.63 in SM with linear regression models fit, illustrating the inverse relationship between presence and reaction time across conditions.
@@ 4.2 Mediation Analysis
We conducted the mediation analysis ( Figure 7) to understand the relationship between interaction, task, and reaction time, with presence as a mediator, by using a bootstrapping approach with 1000 resamples with 95% confidence intervals (CI). The analysis revealed a positive effect of interaction type on presence, with a coefficient of $0.475$ ($p<0.05$, $CI[0.380,0.573]$). This suggests that D resulted in higher levels of presence than S interactions, indicating a strong relationship between interaction type and presence. For the task type, we also examined its effect on presence. The results showed that Select generally resulted in higher levels of presence than Manipulate tasks. For Task Type, the effect on presence was slightly smaller, with a coefficient of $0.362$ ($p<0.05$, $CI[0.245,0.485]$), suggesting that Select tasks led to higher levels of presence than Manipulate tasks. The next phase of our analysis focused on the relationship between presence and reaction time. We found that Presence significantly negatively affected reaction time in the studied conditions, with a coefficient of $-0.610(p<0.05$, $CI[0.245,0.485]$), suggesting that as presence increased, reaction time became faster (also in  Figure 8). However, this effect may be influenced by user familiarity and task type [19], though disentangling these factors is beyond the scope of our study.
@@
The indirect effect of interaction type on reaction time, mediated by Presence, was calculated to be $-0.290(p<0.05$, $CI[-0.398,-0.190]$). This suggests that the improvement in reaction time associated with interaction type was largely mediated by the sense of presence. We also calculated the indirect effect of task type on reaction time, again mediated by Presence. This indirect effect was $-0.221(p<0.05$, $CI[-0.310,-0.130]$), showing that Select improved presence, leading to faster reaction times. The direct effect of interaction type on reaction time, controlling for presence, is found to be $0.017(p>0.05$, $CI[-0.087,0.122]$) 0.017, which is small and not statistically significant. This indicates that once presence is accounted for, interaction type had little direct influence on reaction time. This suggests that the effect of interaction type on reaction time is primarily mediated by presence. The direct effect of task type on reaction time, controlling for presence, is $0.092(p>0.05$, $CI[-0.015,0.199]$). This indicates a small positive effect of Select on reaction time, meaning that this direct effect is not statistically significant. Thus, the mediator explains the relationship between tasks and reaction time.
@@ 4.4 Interpretation of Results
In this section, we interpret the effects identified in our statistical analysis and explore the reasons for their manifestation. Our study used the overall change in presence as the interaction mechanism changed as the first check; the relationship between presence and reaction time would not be observed if the presence did not change. We observed a significant difference in overall presence, which is found to be statistically significant (Table 3). Therefore, we accept H1. We also found that the change in the interaction mechanism changed the overall presence, and the individual subscales showed consistent statistical significance. The interaction mechanisms differed in enabling users to act, the interface quality, how realistic they felt, and the level of involvement they needed to work. As a result, the interaction type is the dominant independent factor that significantly impacted the presence score across subscales. Furthermore, as the direct interactions were more involved than the S, the presence score for DS and DM is higher (Table 2) and statistically significant, and mediation analysis also supports H1.1 as the indirect effect of interaction type on reaction time, mediated by presence, shows that D result in higher presence. Thus, H1.1 can be accepted. The different tasks differed in the level of involvement they expected from the participant; it is expected that changing a task would impact presence. However, we also expected that the task’s nature might impact the interaction type’s effect on presence; for example, the participant may find a given interaction mechanism better suited to one task than the other. However, our results suggested that while task impacts presence, the effect of the two interaction types on presence does not change. We did not observe an interaction effect at the overall questionnaire and subscales levels (omitted for brevity). Therefore, H2 is rejected. Our next hypothesis took a more specific approach to the cross-over between task and interaction, stating that there will be a change in presence if the interaction changes. This is observed, and H2.1 can be accepted.
@@ 4.5 Thematic Analysis
Involvement: As both quantitative and qualitative results suggested, transitioning between different experimental conditions often caused participants to feel a heightened sense of involvement. Comments such as “Being able to control the VE directly was engrossing" and “I felt really part of the virtual world when the interactions were responsive" were common in DM and a few in DS.
Realism: Similarly, the shift to conditions DM and DS was noted by the participants who commented, “..objects looked almost tangible," and "The realistic response of the environment to my actions added to the fun". This suggests that enhancements in visual and interaction realism contribute to the sense of presence.
Interface Quality: Feedback on interface quality often highlighted the limitations of the field of view and control responsiveness. For example, one participant noted, “..small field of view limited my interaction," while another said, “Sometimes the gestures did not register, which pulled me out of the experience" (reported in Select). These comments align with observations where interface limitations negatively impact user experience.
Comfort: The physical comfort of using the headset also played a crucial role in the presence experience. Participants frequently mentioned comfort issues, stating, “The headset felt heavy after a while," or "I experienced some eye strain during the last few trials."
@@ 5 Discussion
Our findings revealed significant interaction and task type effects on presence scores, as indicated by our two-way repeated measures ANOVA. The interaction type showed a higher influence on the presence scores across all subscales, while task type had a more pronounced effect on reaction time. Notably, the interaction between task and interaction type had a negligible impact, suggesting the influence of interaction type on presence and reaction time across different tasks.
@@
Correlation between Presence and Reaction Time. The correlation between presence and reaction time was modestly negative, ranging from -0.42 to -0.63 across different scenarios. Previous studies in immersive VR environments have also observed similar negative correlations between presence and reaction time, suggesting that enhanced presence reduces reaction time. This aligns with our findings, reinforcing the relationship across both D and S interactions within our study. This correlation is visualized in Figure 8, with linear regression lines illustrating the relationship between these measures. The trend indicates that as presence increases, reaction time decreases, suggesting that higher levels of presence in MR can lead to faster responses.
Implication of Task Design. Participants reported higher presence during Select tasks than Manipulate tasks, likely due to the brief, frequent interactions reinforcing a sense of agency and success. Selection tasks provide frequent feedback that aligns with user expectations in MR, enhancing presence. In contrast, Manipulate tasks require sustained motor coordination, potentially shifting focus to task execution. Prior research [66] suggests continuous motor tasks can increase cognitive load and reduce presence. Future work could explore how interaction frequency impacts presence across varied tasks.
Interpretation of Hypotheses.
@@
Our results indicate that participants experienced heightened presence with direct interactions compared to symbolic interactions in the studied tasks. At the same time, this supports H1 and H2; we acknowledge that the outcomes are constrained by the experimental scope and task design. However, this observation also reflects the intuitive nature of direct interactions that closely mimic real-world actions, facilitating a stronger engagement with the VE. H3 was not supported, highlighting that direct interaction’s superior efficacy in enhancing presence is not contingent on the task type. The findings suggest a preference for direct interactions over symbolic ones in MR; however, this preference may vary with different task complexities and interaction goals. H2.1 was confirmed, underscoring that interaction type changes are sufficient to significantly alter the sense of presence. This finding is crucial for designers of MR interfaces, as it emphasizes the importance of choosing appropriate interactions to enhance the overall experience. H3, H3.1, and H3.2 highlight the relationships between interaction types, reaction times, and presence, confirming that interaction types affect the presence and reaction times.
@@
Theoretical and Practical Implications. These results have multifaceted implications. Theoretically, they contribute to understanding how different interactions influence user experience in MR, offering insights into the psychological and physiological aspects of presence. Practically, these findings can guide the development of more effective MR applications, where the choice of interaction type can be tailored to improve both the sense of presence and performance efficiency.
@@ 6 Limitations and Future Work
This section briefly discusses the limitations of our study. We also outline potential future directions for this work.
One challenge we faced was using presence questionnaires originally designed for VR environments in an MR context. Due to the lack of validated MR-specific presence measures, we used VR questionnaires, which offer the most established and reliable way to assess presence. Additionally, the questionnaire score can indicate the level of presence a user is experiencing. However, since reaction time also depends on the task, just looking at it does not provide enough information about the level of presence. At present, reaction time is best suited for measuring the change in presence rather than the absolute level of presence. Further studies should explore whether the effect of task type on reaction time can be modeled, as this may enable the use of reaction time as a standalone absolute measure of presence. In our study, we focused on two interaction types (D and S) and two tasks (Select and Manipulate) for clarity and control, allowing for a focused examination of foundational MR interactions before exploring more complex variations in future research. Our study focused on reaction time as a performance measure due to its real-time reflection of cognitive and motor effort during MR interactions. However, future work could incorporate complementary metrics, such as accuracy, task completion time, and subjective workload, to capture a more comprehensive view of performance, particularly for more complex tasks. Lastly, further research could explore the long-term effects of these interactions on user adaptation and learning. Analyzing prolonged exposure to varied interactions will reveal how users adapt over time, potentially developing new strategies or behaviors. This research could uncover patterns in the correlation between sustained exposure and presence scores and how extended exposure influences reaction times. Exploring multisensory feedback integration can significantly enhance immersion and presence. Researchers could examine how these additional sensory channels influence the relationship between presence and reaction time by incorporating auditory, haptic, and olfactory cues. VE that allow multi-user interactions provide a unique opportunity to study social presence. Researching how users perceive others in shared virtual spaces and how collaboration impacts interactions, overall presence, and reaction time could be an exciting avenue.